# Quantum simulation of the Sachdev–Ye–Kitaev model using time-dependent disorder in optical cavities

### v1

@@ II.1 Single-mode cQED for low-rank SYK
The final effective model expanded in the fermionic single-mode basis $\psi(r)=\sum_{i}\phi_{i}(r)\,c_{i}$ and including normal ordering takes the form
@@ II.2 Trotterized cQED: convergence to SYK
Turning to the two-body couplings, we see that the disorder of the model is solely given by the position-dependent detuning $\Delta_{\rm da}(r)$. As a good approximation, we can assume its mean to be position-independent, so that on average it contributes as a constant to the integral. Thus, we can easily see that the couplings $J_{ik}$ at different fermionic modes (called sites henceforth) have vanishing mean, while their mean is non-zero for equal sites, $\bar{J_{ii}}\neq 0$. This also translates into the $\mathcal{J}^{\rm red.}_{ijij}$ (and permutations), which do not have zero mean. Luckily, all these contribute an overall constant to the Hamiltonian, which we can neglect. On the other hand, for similar reasons, different sets of couplings follow different probability distributions, both in shape and in the variances. For a more thorough explanation, we refer to longer_paper . We divide the couplings into three classes: diagonal, of the form $\mathcal{J}_{ijij}$, Gaussian-distributed and with the largest variance; almost diagonal couplings, of the form $\mathcal{J}_{ijil}$, Gaussian distributed and with a variance similar to the diagonal ones; and off-diagonal, of the form $\mathcal{J}_{ijkl}$, distributed as a $K_{0}$ Bessel function, and with a variance an order of magnitude smaller than the previous ones. All the expressions above should be intended modulo antisymmetric permutations of the indices, and $i\neq j\neq k\neq l$.
In the strict large-$N$ limit, only the off-diagonal couplings contribute, and in that limit the cavity does simulate a cSYK model. However, for intermediate values of $N$, this does not happen, and our numerical simulations of the effective model are affected by these different variances when comparing to vanilla SYK (1). For this purpose, we introduce a modified version (modSYK) that takes this difference into account, such that
$H_{\rm{mSYK}}=\left(\sum_{\{ijkl\}\in D}\mathcal{J}^{D}_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}\right.\\ \;\left.+\sum_{\{ijkl\}\in A}\mathcal{J}^{A}_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}+\sum_{\{ijkl\}\in O}\mathcal{J}^{O}_{ijkl}c^{\dagger}_{i}c^{\dagger}_{j}c_{k}c_{l}\right)\ .$ (8)
Similar to standard SYK, in modSYK all couplings are randomly chosen from independent Gaussian distributions, but $D$, $A$, and $O$ represent the three classes of couplings as motivated from the experimental proposal, each with its own variance. Around $N=10$, diagonal and almost diagonal couplings in the effective model contribute almost equally, but the off-diagonal couplings are negligible, due to the smaller variance. We estimate the cross-over beyond which the off-diagonal elements dominate to be around $N\sim 100$, after which a pure cSYK physics should set in. We have checked scrambling and late-time physics comparing out-of-time-ordered correlators (OTOCs) and the spectral form factor (SFF) between modSYK and the effective microscopic model, finding a good quantitative agreement. For more details, we refer the reader to longer_paper and to Fig. 2.
@@ III.2 Random unitary circuit SYK
Figure 2: a): Average error (14) after Trotterizing the random unitary evolution as a function of $J\Delta t$, for $N=8$ and different total evolution times. b) Average error on the SFF for different values of $N$ and $J\Delta t$, evolving up to $n_{\rm max}=2000$. c) A sample of the SFF for cSYK for $N=10$ (solid line) and the SFF from the trotterized time evolution (13) (bullets). The inset shows a brief period of the evolution at late times.
@@ V Conclusions
Furthermore, this method of simulating SYK appears to be experimentally feasible in current cQED platforms. Control over the light-matter coupling in space and time, close to that introduced in (4), has been demonstrated recently orsi:2024aa . The Trotter-step duration available there is determined by the speed at which independent speckle patterns can be alternated in time. Currently available methods based, e.g., on acousto-optical light modulation liu:2023ab yield timescales of the order of $\mu$s, much slower than $1/\Delta$ but fast enough to allow for thousands of steps within an experimental cycle time. The number of independent patterns that can be created optically is given by the number of independent optical modes available over the atomic cloud. Requiring the number of modes to exceed the number of particles thus amounts to having an optical resolution below the Fermi wavelength. This would be achieved in the cavity-microscope system orsi:2024aa for a typical mesoscopic Fermi gas, such as reported in holten:2022aa . As discussed above, the main limitation will arise from the finite cooperativity, specifying the rate of photon leakage with respect to coherent, virtual photon exchanges Tanji-Suzuki:2011ac . The ultimate limit set by fabrication limitations for optical mirrors suggests that cooperativities of order of a few hundreds could be reached with currently available technologies. Cooperativities on that scale have already been reported at optical wavelengths in microcavities baghdad:2023aa . There, we expect the SYK regime where $N\sim R$ to be reachable using Fermi gases of up to about $50$ atoms, with timescales reaching the scrambling time $\sim\log(N)/J$.

### v2

@@ II.1 Single-mode cQED for low-rank SYK
The final effective model expanded in the fermionic single-mode basis $\psi(r)=\sum_{i}\phi_{i}(r)\,c_{i}\$and including normal ordering takes the form
@@ II.2 Trotterized cQED: convergence to SYK
Turning to the two-body couplings, we see that the disorder of the model is solely given by the position-dependent detuning $\Delta_{\rm da}(r)$. As a good approximation, we can assume its mean to be position-independent, so that on average it contributes as a constant to the integral. Thus, we can easily see that the couplings $J_{ik}$ at different fermionic modes (called sites henceforth) have vanishing mean, while their mean is non-zero for equal sites, $\overline{J_{ii}}\neq 0$. This also translates into the $\mathcal{J}^{\rm red.}_{ijij}$ (and permutations), which do not have zero mean. Luckily, all these contribute an overall constant to the Hamiltonian, which we can neglect. On the other hand, for similar reasons, different sets of couplings follow different probability distributions, both in shape and in the variances. Indeed, one can show that a comparatively large constant mean in $\overline{J_{ii}}$ significantly enhances the variance of diagonal couplings of the form $\mathcal{J}_{ijil}$, and almost diagonal couplings of the form $\mathcal{J}_{ijil}$ (and permutations). By contrast, the only couplings left unaffected are the off-diagonal ones, namely $\mathcal{J}_{ijkl}$ with all indices different, which have the smallest variance and are distributed according to a $K_{0}$ Bessel function. While in the strict large-$N$ limit only the off-diagonal couplings contribute, at intermediate NN–including values realistically attainable in cQED experiments–the variance in the different coupling sectors affects the dynamics of chaos probes. By explicitly separating the constant mean $\overline{J_{ii}}$, it becomes clear that the issue originates from an additional SYK2 contribution that contaminates the chaotic SYK4 dynamics. Fortunately, since this SYK2 is linear in the $J_{ij}$ couplings, it can be dynamically compensated by cyclically switching the detuning $\Delta_{\rm da}$ between positive and negative values, thereby eliminating the corresponding term in the Hamiltonian (such technique is commonly called dipole compensation).
We have then checked standard chaos probes of scrambling and late-time behavior in the resulting model with the constant $\overline{J_{ii}}$ means removed. In particular we have compared out-of-time-ordered correlators (OTOCs) and the spectral form factor (SFF) in this model with those of SYK4 at different values of $R$, finding that the two converge as $R$ increases. For further details, we refer the reader to longer_paper and to Fig. 2.
@@ III.2 Random unitary circuit SYK
Figure 2: a): Average error (13) after Trotterizing the random unitary evolution as a function of $J\Delta t$, for $N=8$ and different total evolution times. b) Average error on the SFF for different values of NN and $J\Delta t$, evolving up to $n_{\rm max}=2000$. c) A sample of the SFF for cSYK for $N=10$ (solid line) and the SFF from the trotterized time evolution (12) (bullets). The inset shows a brief period of the evolution at late times.
@@ V Conclusions
[…]
@@ VI Acknowledgements
SI Supplementary Material
In this Section, we expand on the experimental realization of the effective cavity Hamiltonian, and on dissipation effects. In particular, we give a complete explanation of the various terms that compose the cavity Hamiltonian, and show how they can effectively describe a cSYK4 upon integrating out the cavity photons. Moreover, we analyze the various sources of dissipation that affect the cavity implementation, and show how they compete with the unitary evolution computing the Schwinger–Dyson equations of the system.
SI.1 Effective Cavity Hamiltonian
In this section, we give more information on how to derive the cavity Hamiltonian presented in (4). The microscopic many-body Hamiltonian realised in the envisioned cavity setup is composed of many terms, mentioned above and in Uhrich:2023ddx. In particular, we have
$H=H_{\rm kt}+H_{\rm c}+H_{\rm a}+H_{\rm ac}+H_{\rm ad}\ .$ (S1)
The first contribution, $H_{\rm{kt}}$, describes the collective dynamics of the atoms in the center-of-mass frame. These are represented by the field operators $\psi_{s}(r)$, where $s\in\{g,e\}$, describing the ground and excited state, respectively. With this notation, we have
$H_{\rm kt}=\sum_{s\in\{e,g\}}\int\text{d}^{2}r\,\psi_{s}^{\dagger}(r)\left(\frac{p^{2}}{2m_{\rm at}}+V_{t}(r)\right)\psi_{s}(r)\ .$ (S2)
The integral is two dimensional since we are neglecting any movement on the direction of the cavity axis and runs over the plane of the atomic cloud. In the formula above, $m_{\rm at}$ is the atomic mass (we will use specific values for 6Li atoms), and $V_{t}(r)$ represents the harmonic trapping potential. The second term, $H_{\rm c}$, is the contribution of transverse cavity modes (TCM), written in the rotating frame (see, e.g., Mivehvar_2021 for a recent review of cQED with quantum gases) of the drive $\omega_{\rm d}$,
$H_{\rm c}=\sum_{m}(\omega_{m}-\omega_{\rm d})a_{m}^{\dagger}a_{m}.$ (S3)
The frequency of the different TCM’s is $\omega_{m}=\omega_{0}+m\,\delta\omega$, and in the limit $\delta\omega/\omega_{0}\gg 1$, we can neglect all higher modes with $m>0$. This regime is within the capabilities of existing laboratory setups Brantut_rand_spin_model_2023. The operators $a_{m},a_{m}^{\dagger}$ annihilate respectively create a cavity photon with mode frequency $\omega_{m}$. For further convenience, we also define $\Delta\equiv\omega_{0}-\omega_{\rm d}$. Next, the term $H_{\rm a}$ represents the contribution of the excited state energy of the atoms,
$H_{\rm a}=\int\text{d}^{2}r\,\big(\omega_{\rm a}-\omega_{\rm d}\big)\,\psi_{e}^{\dagger}(r)\psi_{e}(r)\ .$ (S4)
The frequency $\omega_{\rm a}$ corresponds to the transition frequency from ground $|g\rangle$ to excited state $|e\rangle$. We define the detuning $\Delta_{\rm da}\equiv\omega_{\rm a}-\omega_{\rm d}$, and we assume it to be the dominant energy-scale, in order to adiabatically eliminate excited states, leaving the ground state of the $N$ atoms to represent $N$ spinless fermions of the target model, $c_{i}^{\dagger},c_{i}$. This detuning is actually position dependent, $\Delta_{\rm da}(r)$, due to the position-dependent AC Stark shift induced by the speckle pattern. Finally, the last two terms (written in the rotating frame approximation) are
$H_{\rm ac}=\frac{\Omega}{2}\int\text{d}^{2}r\,\left(g_{0}(r)\psi_{e}^{\dagger}(r)\psi_{g}(r)\,a+h.c.\right)\ ,\qquad\qquad H_{\rm ad}=\Omega_{\rm d}\int\text{d}^{2}r\,\left(g_{\rm d}(r)\psi_{e}^{\dagger}(r)\psi_{g}(r)+h.c.\right)\ ,$ (S5)
and represent the interactions between the atoms and cavity, respectively between the atoms and drive modes. These are the QED interaction vertices between charged particles and light, suitably divided between the interaction with the classical field (the drive, from $H_{\rm ad}$), and the quantum field, represented by photons in the lowest mode. $\Omega$ and $\Omega_{\rm d}$ are the coupling strengths between the atoms and the cavity mode, and the atoms and the drive, respectively. Further, $g_{0}(r)$ and $g_{d}(r)$ are the amplitudes of the cavity and drive modes, which in the two-dimensional plane of the atomic cloud are given by Hermite–Gauss modes. In the main text, we set $g_{\rm d}(r)$ and $g_{0}(r)$ to unity, the former in a long-wavelength approximation and the latter so that all $J_{ii}$ couplings have (approximately) the same mean. More details can be found in longer_paper.
The position-dependent adiabatic elimination of excited states is used to induce disorder in the system, as can be seen in (4). In order to imprint it, we consider a third auxiliary state $|\alpha\rangle$, which is a state further excited from $|e\rangle$ by the introduction of the light-shift beam of Rabi frequency $\Omega_{\rm b}$. As such, this light-shift beam causes an AC-Stark shift of the energy of the excited state $|e\rangle$, which in turn induces a shift of the detuning frequency $\Delta_{\rm da}$. Allowing this light-shift to be modulated spatially results in a situation where the effective detuning is now a spatially varying function $\Delta_{\rm da}({r})$, resulting in a disordered structure that filters down via the expansion $\psi(r)=\sum_{i}\phi_{i}(r)\,c_{i}\$to the two-body interaction. This results in the effective Hamiltonian (6). For more detail, see also the multi-mode proposal Uhrich:2023ddx, as well as the companion paper longer_paper.
SI.2 Dissipation
The two main sources of non-Hermitian dynamics in the cavity are two, namely physical photons leaving the cavity, and spontaneous decay from the excited states. These are encoded in the two Hamiltonians of (S5), and they represent corrections to integrating out the photons, and to the adiabatic elimination of the excited states. Generally, in a Markovian approximation, the Lindblad jump operator entering the Liouville dynamics is given by the interaction vertex between the system and the environment. In our case, these are Uhrich:2023ddx
$L_{\Gamma}(r)=$ $\,\frac{\sqrt{\Gamma}\,\Omega_{\rm d}}{\Delta_{\rm da}(r)+i\Gamma/2}\,g_{\rm d}(r)\psi_{\rm g}^{\dagger}(r)\psi_{\rm g}(r)$ (S6)
$L_{\kappa}(r)=$ $\,\frac{\sqrt{\kappa}\,\Omega_{\rm d}\Omega}{2(\Delta-i\kappa/2)}\int\text{d}^{2}r\,\frac{g_{\rm d}(r)g_{0}(r)}{\Delta_{\rm da}(r)+i\Gamma/2}\,\psi_{\rm g}^{\dagger}(r)\psi_{\rm g}(r)$ (S7)
the former modeling photon scattering and the latter photon losses. We then notice that the Lindblad jump operators are quadratic in the fermionic sites, and they depend on the speckle pattern that also generates the $J_{ij}$’s. Therefore, in our case, the jump operators can schematically be written as
$L=\sum_{ij}K_{ij}c_{i}^{\dagger}c_{j}\ ,\qquad\qquad\overline{|K_{ij}|^{2}}=\frac{K^{2}}{N^{2}}\ ,$ (S8)
where the $N$-scaling is set to be equal to the $J_{ij}$, for the aforementioned reason. For a more precise explanation we refer the reader to Uhrich:2023ddx; longer_paper. In both cases, the Lindblad operators arise because of light–matter interaction, and therefore depend on the speckle pattern so that $K_{ij}$ is correlated with $J_{ij}$, and thus such Lindblad couplings are correlated with the cSYK couplings $\mathcal{J}^{\rm red.}_{ijkl}.$ However, as we mentioned in the main text, in the large $R$ limit, they effectively become uncorrelated up to $1/R$ corrections.
Neglecting for simplicity this correction, we can take inspiration from Ryu_2022 to study how this non-Hermitian dynamics affects the system. For simplicity, we restrict to infinite temperature, which is a fixed-point of the Lindblad evolution. Then, in terms of coherent state path integrals, the Lindblad evolution of the density matrix can be rewritten in terms of a Keldysh effective action Sieberer_2016 as [^2] See also this set of lecture notes.
$Z=\int\mathcal{D}[\psi_{\pm},\psi_{\pm}^{\dagger}]e^{iS[\psi_{\pm},\psi_{\pm}^{\dagger}]}\ ,\qquad\qquad S[\psi_{\pm},\psi_{\pm}^{\dagger}]=\int\text{d}t\bigg[\psi_{+}^{\dagger}i\partial_{t}\psi_{+}-\psi_{-}^{\dagger}i\partial_{t}\psi_{-}-H_{+}[\psi_{+},\psi_{+}^{\dagger}]+H_{-}[\psi_{-},\psi_{-}^{\dagger}]\\ -i\sum_{\alpha}\left(2L_{\alpha,-}^{\dagger}L_{\alpha,+}-L_{\alpha,+}^{\dagger}L_{\alpha,+}-L_{\alpha,-}^{\dagger}L_{\alpha,-}\right)\bigg]\ .$ (S9)
The effective action above is quadratic in the jump operators. In order to linearize it, we can employ complex Hubbard–Stratonovich auxiliary fields $\theta_{\alpha,\pm}$ and $\bar{\theta}_{\alpha,\pm}$. The result is
$S[\psi_{\pm},\psi_{\pm}^{\dagger},\theta_{\alpha,\pm},\bar{\theta}_{\alpha,\pm}]=\\ \int\text{d}t\left(\psi_{+}^{\dagger}i\partial_{t}\psi_{+}-\psi_{-}^{\dagger}i\partial_{t}\psi_{-}-H_{+}[\psi_{+},\psi_{+}^{\dagger}]+H_{-}[\psi_{-},\psi_{-}^{\dagger}]\right)+i\int\text{d}t\sum_{\alpha}\left[\begin{matrix}\bar{\theta}_{\alpha,+}(t)\\ \bar{\theta}_{\alpha,-}(t)\end{matrix}\right]^{\mathsf{T}}\left[\begin{matrix}1&0\\ 2&1\end{matrix}\right]\left[\begin{matrix}\theta_{\alpha,+}(t)\\ \theta_{\alpha,-}(t)\end{matrix}\right]\\ -i\int\text{d}t\sum_{\alpha}\Big(\bar{\theta}_{\alpha,+}(t)L_{\alpha,+}(t)+\bar{\theta}_{\alpha,-}(t)L_{\alpha,-}(t)+L_{\alpha,+}^{\dagger}(t)\theta_{\alpha,+}(t)+L_{\alpha,-}^{\dagger}(t)\theta_{\alpha,-}(t)\Big)\ .$ (S10)
It is not hard to verify that integrating out the auxiliary fields we obtain the original Lindblad evolution. We can now simply derive the Schwinger–Dyson equations for the two-point function $G(t)=\langle c_{i}^{\dagger}(t)c_{i}(0)\rangle$ and the self-energy $\Sigma(t)$, as well as for the corresponding ones of the auxiliary fields $G_{ab}^{\theta}(t)=\langle\bar{\theta}_{\alpha,a}(t)\theta_{\alpha,b}(0)\rangle$ and $\Sigma^{\theta}_{ab}(t)$. The Schwinger-Dyson equations for the auxiliary fields are
$\Sigma_{ab}^{\theta}(t)=-\frac{K^{2}}{4}G_{ab}^{2}(t)\ ,\qquad\qquad\mathbf{G}^{\theta}(t)=\left[\begin{pmatrix}1&0\\ -2&1\\ \end{pmatrix}\delta(t)-\mathbf{\Sigma}^{\theta}(t)\right]^{-1}\ ,$ (S11)
while the ones for the fermionic fields are
$\Sigma_{ab}(t)=\,-\frac{J^{2}R}{N}\,s_{ab}G_{ab}^{3}(t)+\frac{K^{2}R}{N}\left(G_{ab}^{\theta}(t)+G_{ba}^{\theta}(-t)\right)G_{ab}(t)\ ,\qquad\qquad\mathbf{G}(t)=\left[\mathbf{G}_{0}^{-1}(t)-\mathbf{\Sigma}(t)\right]^{-1}\ .$ (S12)
The subscripts $a$ and $b$ take values in the set $\{+,-\}$ with the matrix $s_{ab}$ having components $s_{++}=s_{--}=1$ and $s_{+-}=s_{-+}=-1$. The inverses refer to both the $\{a,b\}$ indices, as well as for the time-domain.
Equation (S12) makes it apparent that dissipation contributes at the same order as the unitary evolution in the large-$N$ limit. Given that chaotic timescales are parametrically large333for instance, the scrambling time is $t_{\rm scr}\sim\log(N)$., this implies that in the large-$N$ limit such chaotic dynamics eventually have to compete with the dissipation scale, and their experimental observability depends on the quality factor of the cavity at a given atom number.